OpenStudy (blaque678):

Suppose the graph of f is given. Describe how the graph of the function can be obtained from the graph of f. y=f(-x)

OpenStudy (anonymous):

Dude you could of had went on google and looked this up, I did. Found an answer, hope this helped :) a. y = f(x+5) Shift the graph of y=f(x) to the left 5 units y=f(x+5+5)=f(x+10 Adding 5 makes the x vales occur earlier, i.e. what would happen at x=5 happens at x=0. b. y = f(x) + 5 Shift the graph of y=f(x) upward 5 units. Moving upward means you have to add to the y value. So, instead of y=f(x)+5 it is f(x)+5 +5 or f(x)+10